# Synthetic Data Generation in Cybersecurity: A Comparative Analysis

### v1

@@ V Evaluation Metrics
    Another crucial aspect is the preservation of the data structure. For instance, if a variable is binary in the real data, it must remain binary in the synthetic data to ensure logical consistency. Similarly, suppose a variable is defined to have values greater than or equal to zero (such as packet size in network traffic data). In that case, the corresponding variable in the synthetic data should also respect this constraint. Failing to maintain these structural properties can lead to synthetic data that, while statistically similar on a superficial level, fails to capture the logical and contextual nuances necessary for accurate downstream analyses and applications [108]. In this study, the following metrics were used to measure the statistical similarity (fidelity), as shown in Tables V to X:
    1. (a)
      Data Boundaries (DB): This metric checks whether the synthetic data adheres to the original data’s logical minimum and maximum values. For example, a binary column should contain only binary values, ensuring no out-of-bound entries.
    2. (b)
      Correlation (Corr): This is assessed by examining the correlation heatmap of all variables’ absolute values of the correlation coefficients. It evaluates the strength and direction of relationships in both real and synthetic datasets.
    3. (c)
  2. 2.
    Thus, in this study, the accuracy of TRTR was compared with TSTR for measuring ML utility, as reported in Tables V to X.
  3. 3.
    Ensuring that the class distribution closely mirrors the real data is essential when assessing synthetic data. This involves not only comparing the overall proportions of each class but also examining the stability of these proportions across different subsets of the data. Some methods and tools for generating synthetic data inherently include mechanisms to evaluate and maintain class balance, ensuring that the synthetic dataset accurately reflects the class distribution of the real dataset. Suppose a synthetic data generation method does not inherently address class balance. In that case, it may require additional post-processing to adjust class distributions and avoid introducing bias into the model training process [114].
    Thus, the CB results mentioned in Tables V to X represent how much difference exists between the NORMAL and ATTACK cases. If the classes in the synthetic data are the same, there is zero difference; otherwise, the difference is detailed in the results table later in this paper.
VI Experiment Setup
TABLE I: Summary of 26 Encoded selected Features from NSL-KDD Dataset
@@ VI Experiment Setup
TABLE II: Summary of 21 selected Features from CICIDS-17 Dataset
@@ 20 Flow Bytes/s Rate of flow in bytes per second
TABLE III: Mutual Information Score of NSL-KDD Features
@@
TABLE IV: Mutual Information Score of CICIDS-17 Features
@@ VI-B Software Setup
VI-C Datasets & Pre-processing
  1. 1.
    The preprocessing of the NSL-KDD dataset involved several essential steps to prepare it for synthetic data generation and evaluation. Initially, categorical columns were encoded to transform non-numeric data into a suitable format for ML algorithms. This encoding process ensured that all categorical variables were converted into numerical values, facilitating the subsequent analysis. Following the encoding, feature extraction was performed to identify the most significant features using Mutual Information. This process selected the features’ top 25% (Q1) based on the mutual information weights. This step resulted in a reduced feature set of 26 encoded features (Table I), which were the most important (based on the information weights; Table III) for accurately representing the data. This refined feature set was then used in the data generation phase, ensuring that the synthetic data closely mirrored the critical characteristics of the original dataset.
  2. 2.
    CICIDS-17 dataset, part of the Canadian Institute for Cybersecurity’s intrusion detection dataset collection, is widely employed for evaluating cybersecurity systems [18]. It provides a comprehensive network traffic data set with diverse attack types and normal traffic, offering a robust benchmark for performance assessment. CICIDS-17 has one week of captured network traffic data in eight comma-separated files (CSV). These files are generated based on the attack types. There are 2,830,743 instances of all eight files and 79 columns (78 numerical features and one binary target).
    Preprocessing of the CICIDS-17 dataset involved several vital steps to ready it for synthetic data generation and evaluation. First, all the files were individually tested for null values and duplicated columns. In this process, it was found that there were some infinity values in the data. Thus, all the null values, infinity, and duplicated column ”Fwd Header Length” were removed. Secondly, all the cleaned data files were concatenated into a single CSV file for further feature selection. Lastly, mutual information was performed to select the variables’ top 25% (Q1) based on the mutual information weights. This step resulted in a reduced feature set of 21 features (Table II), which were determined to be the most important (based on the information weights; IV) for accurately representing the data. This reduced feature set was then employed in the synthetic data generation phase, ensuring that the synthetic data accurately reflected the essential characteristics of the original dataset.
VII Experimental Results and Discussion
TABLE V: Comparison of Synthetic Data by Standard Methods on NSL-KDD Dataset
Method Description Pros Cons Performance Summary
SS CB Accuracy
@@ VII Experimental Results and Discussion
ROS [116] Randomly duplicates minority class samples Simple, low cost Overfitting risk, no new samples Yes Yes 0% diff 0% diff 0.9998
SMOTE [63] Generates synthetic samples by interpolation Reduces overfitting, better minority representation May create ambiguous samples, high computation No Yes 0% diff 0% diff 0.9999
Adaptive Synthetic Sampling [67] Adjusts sampling based on data distribution Flexible, effective for imbalance Complex tuning, high computation No Yes 3.8% diff 0.14% diff 0.9998
Cluster based oversampling [117] Oversamples within identified clusters Captures local structures Sensitive to clustering algorithm No Yes 0% diff 0% diff 0.9995
Gaussian Mixture Model [75] Uses Gaussian distributions for sampling Captures complex distributions Assumes data distribution No Yes 99.1% diff 99.9% diff 0.5314
TABLE VI: Comparison of Synthetic Data by Standard Methods on CICIDS-17 Dataset
Method Statistical Similarity Class Balance Accuracy (0.9978)
Data Structure Correlation Probability Distribution
ROS Yes Yes 0% diff 0% diff 0.9991
SMOTE Yes Yes 0% diff 0% diff 0.9990
Adaptive Synthetic Sampling Yes Yes 0% diff 0.2% diff 0.9986
Cluster-based oversampling No Yes 0% diff 0% diff 0.9930
Gaussian Mixture Model No No 61.9% diff 46.1% diff 0.4509
Accuracy” = Machine Learning Utility on TRTR vs TSTR, ”TRTR” = Train and Test on Real Data, ”TSTR” = Train on Synthetic and Test on Real Data
TABLE VII: Comparison of Synthetic Data by AI Methods on NSL-KDD Dataset
Method Description Pros Cons Performance Summary
SS CB Accuracy
DS Corr PD
Bayesian Neural Networks (BN) [118] Uses Bayesian inference for uncertainty estimation Quantifies uncertainty, probabilistic dependencies Requires prior knowledge, complex tuning No No 0% diff 17.2% diff 0.9587
Variational Tabular Autoencoders (TVAEs)[119] Neural networks for unsupervised feature learning Learns representations, reduces dimensionality KL divergence issue, risk of overfitting Yes Yes 23.1% diff 26.7% diff 0.9807
Diffusion Models [86] Uses diffusion for data generation Captures temporal dependencies Needs sequential data, high computational cost No No 98.3% diff 34.2% diff 0.51
Conditional Tabular Generative Adversarial Networks (CTGAN) [12] Adversarial training between generator and discriminator High-quality data, models complex distributions Training instability, mode collapse Yes Yes 11.5% diff 5.7% diff 0.9640
SS = Statistical Similarity, DS = Data Structure, Corr = Correlation, PD = Probability Distribution, CB = Class Balance, ”Accuracy” Machine Learning Utility on TRTR vs TSTR, ”TRTR” = Train and Test on Real Data = 0.9995, ”TSTR” = Train on Synthetic and Test on Real Data
TABLE VIII: Comparison of Synthetic Data by AI methods on CICIDS-17 Dataset
Method Statistical Similarity Class Balance Accuracy (0.9978)
Data Structure Correlation Probability Distribution
Bayesian Neural Networks (BN)
Tabular Variational Autoencoder (TVAE) Yes No 1.4% diff 69.4% diff 0.9718
Conditional Tabular GAN (CTGAN) Yes Yes 0.8% diff 5.2% diff 0.9603
Accuracy” Machine Learning Utility on TRTR vs TSTR, ”TRTR” = Train and Test on Real Data, ”TSTR” = Train on Synthetic and Test on Real Data
TABLE IX: Comparison of Synthetic Data by prominent GANs on NSL-KDD Dataset
Method Statistical Similarity Class Balance Accuracy (0.9995)
Data Structure Correlation Probability Distribution
@@
Accuracy” Machine Learning Utility on TRTR vs TSTR, ”TRTR” = Train and Test on Real Data, ”TSTR” = Train on Synthetic and Test on Real Data
TABLE X: Comparison of Synthetic Data by prominent GANs on CICIDS-17 Dataset
Method Statistical Similarity Class Balance Accuracy (0.9978)
Data Structure Correlation Probability Distribution
@@
Accuracy” Machine Learning Utility on TRTR vs TSTR, ”TRTR” = Train and Test on Real Data, ”TSTR” = Train on Synthetic and Test on Real Data
Given the surge in interest in generating synthetic tabular data, the GAN framework has emerged as a seminal approach, attracting substantial research attention. Consequently, this paper delves deeply into evaluating current high-performing GAN models, leveraging recent advancements in this domain. A meticulous analysis of four leading models is presented in Section III, elucidating their methodologies and performance metrics.
VII-A Results
This study evaluated various methods for generating synthetic tabular data, covering both statistical and AI-based approaches.
VII-A1 Statistical Methods
Five widely recognized statistical methods were assessed, as outlined in the survey by [3] (detailed in Section III). The findings indicate the following:
  • ROS emerged as the most effective statistical method for preserving the data distribution and ensuring class balance. It maintained a high degree of statistical similarity to the original dataset.
@@ VII-A1 Statistical Methods
  • SMOTE and Cluster Centroid outperformed ROS by maintaining statistical similarity and enhancing ML utility, achieving a $99\%$ accuracy while preserving class balance.
  • GMM showed strong performance in capturing feature correlations but underperformed in statistical similarity and ML utility, with an accuracy of only $53\%$.
  • The results were consistent across both the synthetic NSL-KDD and CICIDS-17 datasets (Table V and Table VI).
VII-A2 AI Methods
Four prominent AI and generative AI methodologies were evaluated (details in Section III). The results are summarized as follows:
  • Bayesian Networks struggled with maintaining correlation and structure in the NSL-KDD dataset, especially for binary features. Diffusion models had the lowest ML utility accuracy at $51\%$.
  • TVAE and Conditional Tabular CTGAN performed exceptionally well, preserving statistical properties and achieving high ML utility accuracy rates of $98\%$ and $96\%$, respectively (Table VII).
  • Consistent performance was observed with the CICIDS-17 dataset, where TVAE slightly outperformed CTGAN in ML utility.
VII-A3 Prominent GANs
Four leading GAN models were evaluated in depth (detailed in Section III). The results showed:
  • CTGAN and Copula-GAN outperformed other models, with ML utility accuracy scores of $97\%$ and $98\%$, respectively, on the NSL-KDD and CICIDS-17 datasets.
  • CastGAN showed strong performance in handling feature dependencies but struggled to preserve boundary conditions for specific binary columns.
  • GANBLR++, based on Bayes’ network principles, exhibited performance metrics closely aligned with those of individual BN models.
VII-B Discussion
The results from this study provide several key insights into the effectiveness of various synthetic data generation methods:
VII-B1 Statistical Methods
While effective at maintaining data distribution, ROS is limited in generating genuinely synthetic data, making it less suitable for cybersecurity applications. By definition, ROS does not generate novel data points; it is just random sampling. SMOTE and Cluster Centroid, by enhancing ML utility, present better alternatives for applications where both statistical fidelity and predictive accuracy are critical. Both SMOTE and cluster centroids create novel data points by interpolation and average representation, respectively [63, 120]. GMM’s performance highlights its strength in capturing feature correlations and creating novel data points using the fundamental definition of samples from fitted distributions. Still, its lower utility score suggests it may not be ideal for all applications [121, 122].
VII-B2 AI Methods
Bayesian Networks’ struggle with larger datasets, particularly with multi-class features, indicates a need for better handling such data structures in AI models. The potential reason for this struggle comes from the probabilistic graphical model. This graph tries to represent dependency between variables, and in the case of the CIC-IDS large dataset, it fails to create discretization because of low representation learning (memory) [123, 124]. Diffusion models require further refinement to improve their statistical similarity and utility because they are, in general, designed for high-dimensional data with spatial relationships (e.g., images), making it challenging to capture the non-spatial, relational structure of tabular data [85]. The success of TVAE and CTGAN emphasizes the growing importance of generative AI models in synthetic data generation. However, their lack of inherent class imbalance solutions suggests that additional techniques may be required to address this issue comprehensively.
TVAE and CTGAN excel in generating high-quality synthetic tabular data due to their specialized architectures and loss functions. TVAE uses a probabilistic latent space to model complex dependencies, handling continuous and categorical features with different probability distributions [125]. Its KL divergence-based loss regularizes the model, ensuring diverse yet realistic data generation [126]. CTGAN, on the other hand, uses a conditional generator to capture feature relationships and a specific loss function that optimizes for mode-specific normalization, preserving the distributions of both majority and minority categories [127]. The combination of these tailored architectures and well-designed loss functions enables both models to generate synthetic data that closely mirrors the statistical properties of the original dataset, making them highly effective for tabular data generation [12].
VII-B3 Prominent GANs
The superior performance of CTGAN and Copula-GAN reinforces their suitability for generating synthetic tabular data that retains statistical integrity and practical utility. CopulaGAN excels in generating synthetic tabular data because it uses copulas to model the dependencies between variables and their marginal distributions separately [128]. This allows it to capture complex relationships between features more accurately than standard GANs. By decoupling the modeling of individual feature distributions from their joint dependency structure, CopulaGAN ensures that the synthetic data maintains the statistical properties of the original dataset, making it highly effective for tabular data [129]. CastGAN’s innovative architecture of each generator and auxiliary learner for each variable shows promise, though its limitations with binary features suggest areas for future improvement [92]. The effectiveness of GANBLR++ demonstrates the value of integrating Bayesian principles into GAN frameworks, providing a robust foundation for capturing complex probabilistic relationships in the data for both numerical and categorical datasets [90].
VIII Future Work

### v2

@@ IV Evaluation Metrics
    Another crucial aspect is the preservation of the data structure. For instance, if a variable is binary in the real data, it must remain binary in the synthetic data to ensure logical consistency. Similarly, suppose a variable is defined to have values greater than or equal to zero (such as packet size in network traffic data). In that case, the corresponding variable in the synthetic data should also respect this constraint. Failing to maintain these structural properties can lead to synthetic data that, while statistically similar on a superficial level, fails to capture the logical and contextual nuances necessary for accurate downstream analyses and applications [57]. In this study, the following metrics were used to measure the statistical similarity (fidelity), as shown in results Table VI and Table VII:
    Figure 4: Probability Distribution Comparison for NSL KDD.
    IV-A Experiment Setup
    TABLE II: Summary of 26 Encoded selected Features from NSL-KDD Dataset [58]
@@ IV-A Experiment Setup
    TABLE III: Summary of 21 selected Features from CIC-IDS2017 Dataset [12]
@@ 20 Flow Bytes/s Rate of flow in bytes per second
    1. (a)
      Data Structure (DS): This metric checks whether the synthetic data adheres to the original data’s logical minimum and maximum values. For example, a binary column should contain only binary values, ensuring no out-of-bound entries.
      Table VI and Table VII include a column where each value is calculated based on whether all binary and boolean variables adhere to their expected values. This means that binary variables should be either 0 or 1, and boolean variables should be either TRUE or FALSE. If the synthetic data generated by a method does not meet these structural requirements, the result will be marked as NO.
      For example, in Table VI, the first synthetic data set was generated by ROS, and the value in the DS column is YES, indicating that all binary and boolean variables adhered to their expected values. However, the next row, corresponding to SMOTE, shows NO under the DS column because, apart from the target variable in the NSL-KDD data, the binary and boolean variables did not conform to their expected values. This evaluation method helps assess whether each approach maintains the required data boundaries.
    2. (b)
      Correlation (Corr): This assessment is performed by examining the correlation heatmap, which represents the absolute values of the correlation coefficients for all variables (see Figure 2 and Figure 3 for key NSL KDD heatmaps). It evaluates the strength and direction of relationships in both real and synthetic datasets. If there is a significant difference between the absolute correlation heatmap of the synthetic data and the real data, as well as in the heatmap of the absolute correlation differences, the result is marked as NO. Otherwise, it is marked as YES in the Corr column of all of the results tables.
    3. (c)
      To calculate the statistics for the PD comparison, the following steps were undertaken:
      1. i.
        For every variable in the synthetic and real datasets, the underlying PD was estimated using Kernel Density Estimation (KDE).
      2. ii.
        Each variable’s PD was visualized. For clarity, individual figures showing the real and synthetic PDs were generated. In some cases, side-by-side figures were created to provide a direct comparison (Figure 2). Complete PD comparison graphs for both NSL-KDD and CIC-IDS2017 are provided in Appendix A.
      3. iii.
        The alignment of the PD was quantified using the following formula:
[…]
        where “Number of Variables with Different PD indicates the count of variables whose PD differed between the real and synthetic datasets.
      For the NSL-KDD dataset, which contains 26 total variables, the results are summarized in Table VI. Under the PD column, for the ROS method, all variables exhibited identical PDs according to distribution curves in Figure 4, resulting in a PD of $0\%$ difference. Conversely, for the Adaptive Synthetic Sampling method, one variable out of 26 exhibited a different PD (see Appendix A for a complete graphical PD comparison). Using Equation 1, the percentage difference was calculated as:
[…]
  2. 2.
    Thus, in this study, the accuracy of TRTR was compared with TSTR for measuring ML utility, as reported in Table VI and Table VII.
  3. 3.
    Class Balance Difference (%)=|$\begin{split}\text{Class Balance Difference (\%)}=&\left|P_{\text{synthetic},\text{Normal}}-P_{\text{synthetic},\text{Attack}}\right|\\ &\times 100\end{split}$ (3)
    Thus, the CB results mentioned in Table VI and Table VII represent how much difference exists between the NORMAL and ATTACK cases. If the classes in the synthetic data are the same, there is zero difference; otherwise, the difference percentage is calculated based on the equation. This equation 3 gives 0% diff if the normaland attackclasses are equal, i.e., follow a 50-50 ratio in the whole data. Thus, in Table VI, under the CB column for ROS and SMOTE, it is mentioned 0% diff, and for the Adaptive Synthetic Sampling, it is 0.14% diff.
V Experimental Setup
V-A Mutual Informatiaon
V-A1 Comparison with Other Methods
The rationale for selecting MI over other methods lies in its comprehensive mathematical framework. Initially, this study considered two general approaches for choosing the best features based on their dependence on the target feature and their relationships. Correlation was the first and most obvious choice, mathematically defined as the Pearson correlation coefficient (equ: 4). However, correlation only accounts for linear relationships [69]. In real-world scenarios, non-linear relationships often need to be identified and quantified to select the optimal set of features.
$\text{Corr}(X,Y)=\frac{\text{Cov}(X,Y)}{\sigma_{X}\sigma_{Y}}$ (4)
$IG(Y,X)=H(Y)-H(Y|X)$ (5)
$MI(X,Y)=H(X)+H(Y)-H(X,Y)$ (6)
Compared to other AI-based feature selection methods, such as recursive feature elimination or embedded methods based on decision trees [74], MI has the upper hand by being model-agnostic. Many AI-based techniques are tied to specific algorithms or models, which may introduce biases or limit generalizability across different datasets or classifiers. MI, however, evaluates feature importance based on statistical dependencies independent of any specific model, making it a versatile and unbiased approach [75]. This allows MI to provide a more robust selection process, especially in scenarios where feature relationships are complex and nonlinear, a common occurrence in network data.
V-A2 Application of Mutual Information in this Study
The mutual information score of every feature was determined concerning the target, and the top 25% of the features were selected based on the mutual information weights (Table IV & Table V). This way, the computation was less intensive, the information was high, and the results were more interpretable.
V-B Datasets & Pre-processing
  1. 1.
    The preprocessing of the NSL-KDD dataset involved several essential steps to prepare it for synthetic data generation and evaluation. Initially, categorical columns were encoded to transform non-numeric data into a suitable format for ML algorithms. This encoding process ensured that all categorical variables were converted into numerical values, facilitating the subsequent analysis. Following the encoding, feature extraction was performed to identify the most significant features using Mutual Information. This process selected the features’ top 25% (Q1) based on the mutual information weights. This step resulted in a reduced feature set of 26 encoded features (Table II), which were the most important (based on the information weights; Table IV) for accurately representing the data. This refined feature set was then used in the data generation phase, ensuring that the synthetic data closely mirrored the critical characteristics of the original dataset.
  2. 2.
    CIC-IDS2017 dataset, part of the Canadian Institute for Cybersecurity’s intrusion detection dataset collection, is widely employed for evaluating cybersecurity systems [12]. It provides a comprehensive network traffic data set with diverse attack types and normal traffic, offering a robust benchmark for performance assessment. CIC-IDS2017 has one week of captured network traffic data in eight comma-separated files (CSV). These files are generated based on the attack types. There are 2,830,743 instances of all eight files and 79 columns (78 numerical features and one binary target).
    Preprocessing of the CIC-IDS2017 dataset involved several vital steps to ready it for synthetic data generation and evaluation. First, all the files were individually tested for null values and duplicated columns. In this process, it was found that there were some infinity values in the data. Thus, all the null values, infinity, and duplicated column ”Fwd Header Length” were removed. Secondly, all the cleaned data files were concatenated into a single CSV file for further feature selection. Lastly, mutual information was performed to select the variables’ top 25% (Q1) based on the mutual information weights. This step resulted in a reduced feature set of 21 features (Table III), which were determined to be the most important (based on the information weights; V) for accurately representing the data. This reduced feature set was then employed in the synthetic data generation phase, ensuring that the synthetic data accurately reflected the essential characteristics of the original dataset.
TABLE IV: Mutual Information Score of NSL-KDD Features
@@ V-B Datasets & Pre-processing
TABLE V: Mutual Information Score of CIC-IDS2017 Features
@@ V-B2 Software Setup
VI Results and Discussion
TABLE VI: Comparison of Synthetic Data on NSL-KDD Dataset
Category Method SS CB Accuracy
@@ VI Results and Discussion
Non-AI (Statistical) ROS Yes Yes 0% diff 0% diff 0.9998
SMOTE No Yes 0% diff 0% diff 0.9999
ADASYN No Yes 3.8% diff 0.14% diff 0.9998
CC No Yes 0% diff 0% diff 0.9995
GMM No Yes 99.1% diff 99.9% diff 0.5314
AI-Based (Classical + Generative) BN No No 0% diff 17.2% diff 0.9587
TVAE Yes Yes 23.1% diff 26.7% diff 0.9807
TABDDPM No No 98.3% diff 34.2% diff 0.51
@@
TABLE VII: Comparison of Synthetic Data on CIC-IDS2017 Dataset
Category Method SS CB Accuracy
DS Corr PD
Non-AI (Statistical) ROS Yes Yes 0% diff 0% diff 0.9991
SMOTE Yes Yes 0% diff 0% diff 0.9990
ADASYN Yes Yes 0% diff 0.2% diff 0.9986
CC No Yes 0% diff 0% diff 0.9930
GMM No No 61.9% diff 46.1% diff 0.4509
AI-Based (Classical + Generative) BN
TVAE Yes No 1.4% diff 69.4% diff 0.9718
TABDDPM
@@
This section presents an in-depth comparative evaluation of twelve synthetic data generation methods applied to the NSL-KDD and CIC-IDS2017 datasets. The analysis is structured around four key objectives: fidelity, utility, class balance, and computational efficiency. Tables VI and VII summarize these results for both datasets.
The evaluation focuses on two distinct categories of methods: Non-AI (Statistical) and AI-based (Classical + Generative AI). The performance of these methods is assessed based on statistical similarity (SS), class balance (CB), and machine learning utility (measured as accuracy in TRTR vs. TSTR). Additionally, AI-based methods’ computational cost and training efficiency are critically analyzed.
VI-A Statistical Methods
Statistical methods such as ROS, SMOTE, ADASYN, and Cluster Centroids (CC) are widely used for addressing class imbalance in network traffic datasets. However, these methods primarily rely on resampling strategies rather than truly modeling the underlying data distribution. While SMOTE and ADASYN generate synthetic samples, they do so through interpolation, which may fail to capture complex feature dependencies. Additionally, statistical oversampling techniques can lead to overfitting in downstream machine learning tasks, as the generated samples lack inherent variability compared to real network traffic. Prior studies [X, Y] indicate that such methods are effective in reducing bias in classification but perform poorly when high-dimensional feature interactions are critical for network anomaly detection. They exhibit excellent class balance maintenance, achieving 0% CB difference in most cases. However, these methods primarily resample existing data points rather than generating truly novel synthetic data, which limits their effectiveness in capturing complex relationships between features.
@@ VI-A Statistical Methods
  • SMOTE and CC outperform ROS by enhancing ML utility while preserving statistical similarity.
  • GMM performs poorly due to its parametric limitations, struggling to model high-dimensional network traffic distributions (99.1% PD diff in NSL-KDD, 61.9% in CIC-IDS2017).
  • Despite their fidelity limitations, statistical methods achieve near-perfect accuracy (0.999) in TRTR and TSTR settings, showing they can still be useful for ML training in balanced data scenarios.
VI-B AI-Based Methods
AI-based methods exhibit stronger generative capabilities but face trade-offs in fidelity, utility, and computational cost:
  • CTGAN and CopulaGAN achieve the best balance, with high fidelity (Corr: Yes, PD diff: 7.7% (NSL-KDD), Corr: Yes, PD diff: 0% (CIC-IDS2017)) and strong utility (accuracy: 0.9667 (NSL-KDD) and 0.9538 (CIC-IDS2017)).
  • TVAE is overall highly effective in NSL-KDD but struggles with class balance in CIC-IDS2017 (CB diff: 69.4%), indicating possible limitations in handling imbalanced network traffic distributions.
  • Diffusion models (TABDDPM) perform poorly, particularly in high-dimensional tabular data, due to their reliance on iterative denoising rather than structured feature learning (PD diff: 98.3% (NSL-KDD)).
VI-C Computational Cost and Training Efficiency
The computational efficiency of AI-based methods varies significantly, impacting their scalability for large-scale network traffic simulation:
  • Bayesian Networks (BN) failed on CIC-IDS2017 (Table VII) due to excessive memory requirements when constructing initial feature relationships. This highlights the challenge of applying probabilistic graphical models to large network traffic datasets.
  • TABDDPM could not generate results for CIC-IDS2017 (Table VII), indicating prohibitively high computational costs. Diffusion-based models, while powerful for image data, struggle with tabular datasets due to their need for extensive forward and reverse sampling processes.
  • CTGAN and CopulaGAN balance computational efficiency and fidelity, making them the most practical for synthetic network traffic data where scalability is critical.
Key technical inference from the abovementioned results is that GAN-based models, such as CTGAN and CopulaGAN, perform exceptionally well because they utilize mode-specific normalization and conditional generation, which helps preserve the feature dependencies crucial for generating realistic network traffic samples. This ability to maintain the relationships between features contributes to their effectiveness in generating high-quality synthetic data [32, 76].
On the other hand, diffusion models face significant challenges, particularly when applied to large tabular datasets. These models rely on iterative denoising, a process that becomes computationally expensive as the dataset size increases. Unlike image data, which has spatial dependencies that aid the diffusion process, tabular data lacks this inherent structure, making the data generation process less efficient. Furthermore, diffusion models struggle with discrete categorical features, which are common in network traffic datasets, further hindering their performance [77].
Gaussian Mixture Models also fail to capture the complex nature of network traffic data. These models assume that the data follows underlying Gaussian distributions, but they are unable to account for the multimodal and varied traffic behaviors typically found in network traffic datasets [78].
For practical recommendations, techniques like SMOTE and ROS are still viable options for addressing class imbalance in network traffic datasets. When it comes to high-fidelity data generation, CTGAN and CopulaGAN are the top performers, surpassing other models in generating realistic and useful synthetic data. However, Bayesian Networks and Diffusion Models should be avoided for large-scale datasets due to their high computational costs and inefficiencies in handling such data.
VII Future Work